1. The weight of an astronaut plus his space suit on the Moon is only 350 N. (a) How much does the suited astronaut weigh on Earth? (b) What is the mass on the Moon? On Earth?
  2. A rocket sled accelerates at a rate of 59.0 m/s2 . Its passenger has a mass of 85.0 kg. (a) Calculate the horizontal component of the force the seat exerts against his body. Compare this with his weight using a ratio. (b) Calculate the direction and magnitude of the total force the seat exerts against his body.
  3. Suppose the ski patrol lowers a rescue sled and victim, having a total mass of 80.0 kg, down a 60.0° slope at constant speed, as shown below. The coefficient of friction between the sled and the snow is 0.100. (a) How much work is done by friction as the sled moves 20.0 m along the hill? (b) How much work is done by the rope on the sled in this distance? (c) What is the work done by the gravitational force on the sled? (d) What is the total work done?
1
Expert's answer
2021-08-10T08:00:58-0400

As we know, the acceleration due to gravity is 1/6 of that on earth. So, the weight on the Moon is


"W_M=m\frac g6,\\\space\\
m=\frac{6W_M}{g}=\frac{6·350}{9.8}=214\text{ kg}."

The mass is defined by the volume and density of materials, so, the mass on the Moon is the same as the mass on the Earth.